Find the general solution of each differential equation, expressing y explicitly in terms of x.

1) dy/dx = 2y

2) dy/dx = y+1 / x+2


1
Expert's answer
2021-09-14T06:05:35-0400

1)

"dy\/dx=2y"

"\dfrac{dy}{y}=2dx"

Integrate

"\int\dfrac{dy}{y}=\int2dx"

"\ln|y|=2x+\ln C"

"y=Ce^{2x}"

2)

"\dfrac{dy}{dx}=\dfrac{y+1}{x+2}"

"\dfrac{dy}{y+1}=\dfrac{dx}{x+2}"

Integrate

"\int\dfrac{dy}{y+1}=\int\dfrac{dx}{x+2}"

"\ln|y+1|=\ln|x+2|+\ln C"

"y+1=C(x+2)"

"y=Cx+2C-1"